OpenStudy (anonymous):

The point A(-3, -2) is translated using T: (x,y) → (x + 5, y - 3). What is the distance from A to A'?

OpenStudy (precal):

x coordinate is moved to the right 5 units that is why it is x+5 the y coordinate is moved down 3 units, that is why it is y-3

OpenStudy (anonymous):

because its moving down

OpenStudy (precal):

point is moved 5 units to the right and 3 down

OpenStudy (anonymous):

A' coordinates are (2,-5).Distance between A and A' is \[\sqrt{((2-(-3))^2+((-5-(-2))^2}=\sqrt{34}\]

OpenStudy (anonymous):

Why not A) calculate A and A', then apply distance formula or B) apply distance formula to two general points, then plug in for \(x\) and \(y\)?

OpenStudy (anonymous):

@burakb, we do not give full answers here. See the Code of Conduct.